OpenStudy (anonymous):

what is the third order derivative of (x^-(1/2x))/(x^x) in algebraic terms?

OpenStudy (anonymous):

\[x ^{-1\div2x}\div x^{x}\]

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

ellipsis

OpenStudy (anonymous):

ellipsis

OpenStudy (king):

hashir????

OpenStudy (anonymous):

hashir????

OpenStudy (anonymous):

the first one \[y'=-1.5*x ^{-1.5x}*(lnx+1)\]

OpenStudy (anonymous):

bro i am jst enjoying !!

OpenStudy (anonymous):

in non-logistical terms?

OpenStudy (anonymous):

in non-logistical terms?

OpenStudy (lalaly):

first derivative\[\large{\frac{x^{-\frac{1}{2}x}}{x^x}= x^{-\frac{3}{2}x}}\]let \[y=x^{-\frac{3}{2}x}\]\[lny=\ln(x^{-\frac{3}{2}x})\]\[lny=-\frac{3}{2}xlnx\]now take derivative of both sides\[\frac{1}{y}y'=-\frac{3}{2}lnx-\frac{3}{2}\]\[y'=(\frac{-3}{2}lnx-\frac{3}{2})y\]

OpenStudy (anonymous):

first derivative x−12xxx=x−32x let y=x−32x lny=ln(x−32x) lny=−32xlnx now take derivative of both sides 1yy′=−32lnx−32 y′=(−32lnx−32)y

OpenStudy (lalaly):

Hashir stop this please

OpenStudy (lalaly):

He left:D

OpenStudy (anonymous):

thee second \[y''=1.5x ^{-1.5x}*(1/x+1.5(lnx+1)^2)\]